in a given city 6% of all drivers get at least one parking ticket per year . use the position approximation to the binomial distribution to determine the probabilities that among 80 drivers (randomly choosen in the city). (1) 4 will get at least one parking ticket in any given year, (2)at least 3 will get one parking ticket in any given year, (3)anywhere from 3 to 6 inclusive ,will get at least one parking ticket in any given year​

1
Expert's answer
2022-06-08T15:59:05-0400

"n=80>20, np=80(0.06)=4.8<5"

The Poisson distribution approximates the binomial distribution


"\lambda=np=4.8"

(1)


"P(X=4)=\dfrac{e^{-4.8}(4.8)^4}{4!}=0.182029"

(2)


"P(X\ge3)=1-P(X=0)-P(X=1)"

"-P(X=2)=1-\dfrac{e^{-4.8}(4.8)^0}{0!}"

"-\dfrac{e^{-4.8}(4.8)^1}{1!}-\dfrac{e^{-4.8}(4.8)^2}{2!}"

"=0.857461"

(3)


"P(3\le X\le6)=P(X=3)+P(X=4)"

"+P(X=5)+P(X=6)"

"=\dfrac{e^{-4.8}(4.8)^3}{3!}+\dfrac{e^{-4.8}(4.8)^4}{4!}"

"+\dfrac{e^{-4.8}(4.8)^5}{5!}+\dfrac{e^{-4.8}(4.8)^6}{6!}=0.648265"